# Epidemiology of Large Language Models: A Benchmark for Observational Distribution Knowledge

### v1

@@ 2.3 Axis 3: Models
Closed‐weight models are available only through API‐based access to proprietary weights, and include a state-of-the-art reasoning model. In particular, we evaluate Reasoning Language Models (RLMs) such as OpenAI’s o4-mini. We also include models with web access, such as OpenAI’s GPT-4.1. Appendix LABEL:appendix:closed-models analyzes the performance of selected closed models.
@@ 2.4 Benchmark Construction
Our benchmark is designed with reproducibility and extensibility as key principles. All dataset constructions are made available, ensuring transparency and verifiability. The modular structure of the benchmark allows easy addition of both new datasets and new models, facilitating broader future evaluations. Furthermore, model evaluation is streamlined: any model available through Hugging Face (wolf2019huggingface) can be benchmarked with a single line of code, allowing for rapid inclusion of diverse architectures. At the time of submission, our benchmark consists of 10 pre-processed datasets, 169 tasks (75 low-dimensional and 94 high-dimensional), with 12 open-weight models evaluated.
3 Key Insights and Observations
Results summarizing the leaderboards for the low- and high-dimensional settings are shown in Fig. 3(b). They indicate that the observational distributions $\tilde{P}(V)$ encoded in the LLMs are closer to the ground truth, real-world observational distributions $P(V)$ than the uniform distribution $P^{\text{unif}}(V)$ since models achieve scores above zero. All models achieve scores better than uniform guessing, suggesting some access to L1 knowledge. However, overall, the models’ performance is rather poor, with the best models scoring an average of 22/100 points in the low-dimensional, and 17/100 in the high-dimensional setting, respectively. Therefore, capabilities of LLMs for encoding real-world observational distributions seem limited, and using such knowledge for downstream tasks should be done with care. Furthermore, Appendix LABEL:appendix:fine-tune shows that fine-tuning may offer limited added benefit, raising interesting questions on how to design approaches for improving the models’ L1 knowledge.
(a) Performance by dataset (low-dimensional).
(b) Performance by dimension.
Figure 4: Model performance across (a) datasets; (b) dimension.
@@ 3 Key Insights and Observations
The performance of models on 75 low-dimensional tasks, comparing $\tilde{P}(V_{Y}\mid V_{X})$ with the ground truth $P(V_{Y}\mid V_{X})$ for $|V_{Y}|=|V_{X}|=1$ is shown in Fig. 3(b) (purple bars). Among a selection of models from different organizations (Meta, Microsoft, Google, DeepSeek AI, Mistral AI), the overall performance across models does not exhibit large differences, and all models perform poorly. Another interesting aspect is the performance of models across different datasets, visualized in Fig. 4(a). The best performance is observed over the FBI Arrest Statistics (sex/race distributions over crime types), ACS Census data (questions on employment, education, and income across demographics), and Departments of Education (sex by graduation degree) and Labor (sex/race by occupation). For three of these four datasets (FBI, BLS, IPEDS) the detailed statistics we queried are available on their websites, meaning that models could have had access to the exact probability tables used for constructing our questions. Therefore, improved performance may correspond to the actual data being available in the model’s training set. This observation also emphasizes the importance of the high-dimensional setting – in which we condition on multiple observed variables – as this knowledge is not readily available on the web.
@@
The performance of models in the high-dimensional case, for $P(V_{Y}\mid V_{X})$ with $|V_{Y}|=1,|V_{X}|\in\{2,3,4,5\}$, and $V_{Y}$ binary, is shown in Fig. 3(b) (yellow bars). As the scores indicate, performance is again quite poor across models, and certainly worse than in the low-dimensional case. We also inspect the model performance with varying dimension (Fig. 4(b)). The pattern of the performances does not clearly indicate the curse of dimensionality, possibly due to the fact that each dimension $d$ has a relatively small number of tasks, and additional tasks may need to be added to investigate this effect in more detail.
@@
Recovery of high-dimensional conditional distributions investigated in this work may be seen as related to questions of causal inference. Commonly, to adjust for confounding by a set of covariates $V_{Z}$, data analysts compute the propensity of a treatment $V_{X}$, that is, reweigh the samples with $P(V_{X}=v_{X}\mid V_{Z}=v_{Z})^{-1}$ to obtain estimates of a desired potential outcome (rubin1977assignment; rosenbaum1983central). Other strategies use a predictive model for the outcome for the outcome distribution $P(V_{Y}\mid V_{X},V_{Z})$ in order to allow for or improve causal inference (bang2005doubly). However, if such conditional distributions are not available to a model, eliciting them may be a futile exercise. Therefore, our work brings into question some recent approaches claiming to improve causal inferences using foundation models (de2025efficient). In our view, using AI models for causal inference may require a more sophisticated approach.
@@
A limited number of models are evaluated on our benchmark. However, by choosing a representative sample of models, we cover a range of models developed by different organizations. Furthermore, our emphasis on easy access and evaluation of models will facilitate the benchmarking of other models in future work. Another limitation is that high-dimensional inference is investigated across 4 datasets, and the number of contexts for evaluation should grow over time. We further envision that our work will set the basis for evaluating capabilities of AI models for inference in Layers 2 & 3 of the PCH – to systematically test how observational distributions available to models affect downstream tasks such as causal or counterfactual inferences.
Ethics statement.
This work introduces a benchmark based on publicly available datasets and models, with all assets cited and used in accordance with their licenses. No new data collection, human subjects, or sensitive information are involved. We do not foresee ethical risks such as privacy violations, bias amplification, or harmful applications beyond those already present in the underlying datasets and models. In fact, our benchmark may contribute positively by encouraging researchers to critically evaluate model capabilities for observational knowledge before deployment in downstream applications such as causal inference.
Reproducibility statement.
Strong emphasis is placed on reproducibility. All datasets used are publicly accessible, and we release preprocessing scripts, evaluation code, and instructions to reproduce every reported result. The benchmark is designed to make reproducing results and evaluating new models straightforward. We provide all experimental details, including data splits, hyperparameters, evaluation metrics, and compute resources. Code and documentation are made available in the code repository https://github.com/dplecko/llm-epidemia.
@@
The source code for reproducing the benchmark can be found in the anonymized code repository https://anonymous.4open.science/r/llm-epidemia-563D. The repository includes a README file explaining all the steps for setting up the benchmark. For the Llama3 8B, Mistral 7B, and Phi4 models the experiments were run on a single NVIDIA H100 GPU. Evaluating both the low- and high-dimensional settings required <1 hour per model. Evaluating DeepSeek 7B (Chat version), Gemma3 27B, and LLama3 70B models was done on a node with four NVIDIA GH200 Grace Hopper chips, and running both settings required <1 hour of compute per model.
@@ 74 SCF: Debt by Education
Table 2: High-dimensional tasks. The third column indicates the task dimension $d=|V_{X}|$, and the fourth column indicates whether the task is used for the evaluation of closed models in Appendix LABEL:appendix:closed-models.
@@ Appendix B Comparison of Base & Instruction-Tuned Models
Appendix C Closed Models Performance
In this appendix, we look at how the performance of closed models compares to the performance of open-weight models. In particular, we evaluate two closed models from OpenAI – GPT-4.1 (achiam2023gpt) and o4-mini (openai2025o4mini), where the latter model is a reasoning language model (RLM) (besta2025reasoninglanguagemodelsblueprint). Since these models require access to an API, we evaluate them on a subset of the tasks in the high-dimensional setting (consisting of tasks with a univariate binary $V_{Y}$, and $|V_{X}|\in\{2,3,4,5\}$). In particular, we select 63 out of 94 tasks (the selected tasks are indicated in the Closed Eval column of Tab. LABEL:tab:high-dim), focusing on tasks for which at most 250 queries need to be evaluated on the model.
To elicit the model’s distribution $\tilde{P}(V_{Y}\mid V_{X})$, we use the likelihood-prompting technique, described in detail in Appendix LABEL:appendix:qa-vs-prob. We choose likelihood prompting, since closed models such as GPT-4.1 and o4-mini do not provide access to next-token prediction probabilities, meaning that question/answer-prompting used in the main text becomes possible only via Monte Carlo sampling. This, however, would increase the number of queries sent to the models and the associated cost by at least two orders of magnitude. This is due to the fact that even with $n_{mc}=100$ Monte Carlo samples, the 95% confidence interval for a probability $\tilde{P}(V_{Y}=1\mid V_{X}=v_{X})\in[0,1]$ is only guaranteed to be reduced to a width of approximately 0.1, still reflecting a high level of uncertainty.
The performance of GPT-4.1 and o4-mini models on the 63 selected tasks is shown in Fig. LABEL:fig:closed-models.
The figure shows that GPT-4.1 and o4-mini outperform the evaluated open-weights models. This may be explained by the fact that GPT-4.1 is the largest model we evaluated, in terms of the parameter count (even though the number of parameters is not officially published). Even though OpenAI’s models outperform the evaluated open-weights models, they still do not perform well. In Fig. LABEL:fig:closed-models, the orange line indicates the performance of the so-called mean-baseline, which outputs the overall population mean $P^{\text{mean}}(V_{Y}=1\mid V_{X}=v_{X})=\mathbbm{E}_{P}[V_{Y}]$ for all values $v_{X}\in\mathrm{dom}(V_{X})$ (and extended discussion on this baseline can be found in Appendix LABEL:appendix:mean-baseline). Therefore, none of the models are able to outperform such a baseline.
Retrieval-Augmented Generation (RAG) (lewis2020retrieval).
Retrieval-augmented generation (RAG) is a technique in which an LLM is combined with an external retriever (lewis2020retrieval). In this case, the model does not have to rely only on its internal knowledge, but relevant documents can be fetched from a collection of texts. To investigate the impact of RAG on our benchmark, we used the GPT 4.1 model with web access (the model can search the web for each query). To test how RAG affects performance, we focused on 11 tasks on which the GPT 4.1 model without web-based RAG had a score of 0 (corresponding to task indices 40, 43, 57, 62, 63, 64, 65, 66, 67, 70, and 71 in Tab. LABEL:tab:high-dim). Here, the expectation is that RAG would help the model performance, and would result in a higher score. However, the results for the GPT 4.1 model with web-based RAG still showed a zero score for each task. Therefore, perhaps surprisingly, for these tasks RAG did not result in improved performance. This finding warrants further investigation of how RAG may improve the performance of models on our benchmark. This investigation is left for future work.
Appendix D Details on Responses and Scoring
As mentioned in Sec. 2.2.1, we use a known strategy for eliciting model responses from previous works on question answering (dominguez2024questioning; hendrycks2020measuring; santurkar2023whose). As models may exhibit ordering bias (dominguez2024questioning), we consider all of the different $|\mathrm{dom}(V_{Y})|!$ permutations of labels $A,B,C,\dots$, over the answers $v_{Y}\in\mathrm{dom}(V_{Y})$, and average the probabilities accordingly (if the number of permutations $|\mathrm{dom}(V_{Y})|!>120$, we consider 120 random permutations). The above way of eliciting responses may be used for all models with open-source weights, which allows us to access next-token prediction probabilities. For the ground truth distribution $P$, in the low-dimensional setting we use a simple bin-counting estimator on the available dataset, whereas in the high-dimensional case we fit the distribution $P(V_{Y}\mid V_{X})$ using lightgbm (ke2017lightgbm), with actual values $P(V_{Y}=v_{Y}\mid V_{X}=v_{x})$ obtained out-of-sample, using 5-fold cross-validation.
Appendix E Likelihood vs. Question-Answering Prompts
In this appendix, we discuss an alternative way of eliciting model distributions $\tilde{P}(V_{Y}\mid V_{X})$. We recall Ex. 1 from the main text, in which we were interested in marijuana usage ($V_{Y}$) over age groups ($V_{X}$). For this task, we used a prompt template
$\pi(v_{X})=\text{``Has a person aged \{$v_{X}$\} ever used marijuana?"},$ (8)
which for different values of age $v_{X}$ creates different questions posed to the model. With the question, the model is also given the choice of answers yes/no, which are labeled with letters A, B. For extracting the model’s probability $\tilde{P}(V_{Y}=1\mid V_{X}=v_{X})$, we investigate the probabilities associated with letters A, B in the model’s next-token prediction distribution $\tilde{P}^{\mathrm{nt}}$.
However, this approach, which we refer to as question/answer-prompting (QA-prompting, for short), is not the only way to elicit the model’s distribution $\tilde{P}(V_{Y}\mid V_{X})$. An alternative way of eliciting this distribution is to use a different prompting strategy, which asks for the probability directly:
[…]
This strategy, which we refer to as likelihood-prompting, is investigated in this appendix. To elicit a probability from the model, a range of responses is offered, namely responses in the following form:
    A. 0%
    B. 0%-5%
    C. 5%-10%
    U. 95%-100%
    V. 100%
We then again inspect the next-token prediction probabilities in $\tilde{P}^{\mathrm{nt}}$, and choose the response with the highest probability:
The chosen response letter $\mathcal{A}^{\text{resp}}$ is then mapped to either an interval, or a fixed value of 0% or 100%. For assigning the final predicted probability, for intervals, we take the midpoint. For instance, if the answer B. 0%-5% is chosen, we set the predicted probability to $0.025$, and so on. In this way, we elicit the probability $\tilde{P}(V_{Y}=1\mid V_{X}=v_{X})$ of the model.
Model performances on the benchmark when using likelihood-prompting (compared with QA-prompting), on the 94 tasks in the high-dimensional setting with a binary one-dimensional $V_{Y}$, and $|V_{X}|\in\{2,3,4,5\}$, are shown in Fig. LABEL:fig:prob-leaderboard. On top of the models used in the main text, we also include a version of DeepSeek R1 (guo2025deepseek) distilled using Qwen2.5 32B (hui2024qwen2).
The results in the figure compare the performance of models on the same tasks, where likelihood-prompting is used instead of QA-prompting used in the main text. We note that the model performances, for some models, improve with likelihood-prompting (in particular, the Phi4 model moves from an average score of 17/100 to 33/100, Gemma3 27B from 13/100 to 33/100, LLama3 70B from 7/100 to 27/100). At the same time, for some models, the performance with likelihood prompting becomes worse (LLama3 8B). However, overall, model performance is still quite poor, and we further contextualize this further in the sequel.
E.1 Baselines – how good are the models?
In the evaluation of our benchmark, we used two baselines, which corresponded to the notion of zero probabilistic knowledge. The first baseline was the uniform baseline $P^{\text{unif}}$, which is in the binary case given by:
$P^{\text{unif}}(V_{Y}=1\mid V_{X}=v_{X})=\frac{1}{2}\;\forall v_{X}.$ (11)
This baseline corresponds to random guessing. In the main text, we also discussed the 0/1 baseline, which outputs either a probability 0 for events with a marginal probability $\leq 0.5$, or a probability $1$ for events with a marginal probability >0.5>0.5:
where $\mathbbm{E}_{P}[V_{Y}]$ is the true mean of $V_{Y}$ (in the ground truth distribution $P$). In our scoring scheme, these baselines were used to determine the score $S=0$ for a given task (see Sec. 2.2.1). However, another interesting baseline is the mean-baseline, defined by:
This baseline outputs a fixed value (the marginal mean of the outcome $V_{Y}$) for each $v_{X}\in\mathrm{dom}(V_{X})$. In words, this baseline is aware of the marginal probability of the event $V_{Y}$, but has no knowledge about the variation of $P(V_{Y}=1\mid V_{X}=v_{X})$ according to $v_{X}$ around the marginal mean $\mathbbm{E}_{P}[V_{Y}]$. This baseline corresponds to some probabilistic knowledge (marginal mean), but without any knowledge over how the probability varies across population subsets. Assigning a score of $S=0$ to this baseline would make high scores on our benchmark even more difficult to achieve, and an argument could be made for including it in future iterations of the benchmark. In Fig. LABEL:fig:prob-leaderboard, the performance of the mean-baseline with respect to the current evaluation is shown (orange horizontal line), indicating an average score of 46/100. This means that the inclusion of this baseline would effectively render the scores of all models investigated in this manuscript close to 0. This observation once again confirms that the current generation of LLMs does not possess knowledge about observational distributions in the real world, regardless of the strategy (QA- or likelihood-prompting) we choose for eliciting the model distributions $\tilde{P}(V_{Y}\mid V_{X})$.
Appendix F Does Fine-Tuning Improve Performance?

### v2

@@ 2.3 Axis 3: Models
Closed‐weight models are accessible only via API calls to proprietary weights. In this category, we evaluate state-of-the-art Reasoning Language Models (RLMs), including OpenAI’s o4-mini. We also include closed models with web-augmented capabilities, which rely on retrieval-augmented generation (RAG) (lewis2020retrieval), such as OpenAI’s GPT-4.1.
@@ 2.4 Benchmark Construction
Our benchmark is designed with reproducibility and extensibility as key principles. All dataset constructions are made available, while the modular structure of the benchmark allows easy addition of both new datasets and new models, facilitating broader future evaluations. Furthermore, model evaluation is streamlined: any model available through Hugging Face (wolf2019huggingface) can be benchmarked with a single line of code (see our HF repository), allowing for rapid inclusion of diverse architectures. At the time of submission, our benchmark consists of 10 pre-processed datasets, 169 tasks (75 low-dimensional and 94 high-dimensional), with 12 open-weight models evaluated.
3 Results
Overall Results.
Results summarizing the leaderboards for the low- and high-dimensional settings are shown in Fig. 4, and results can be viewed interactively on the project webpage. They indicate that the observational distributions $\tilde{P}(V)$ encoded in the LLMs are closer to the ground truth, real-world observational distributions $P(V)$ than the uniform distribution $P^{\text{unif}}(V)$ since models achieve scores above zero, suggesting some access to L1 knowledge. However, overall, the models’ performance is relatively poor. For QA prompting, best models scored an average of 22/100 points in the low-dimensional, and 17/100 in the high-dimensional setting, while for likelihood prompting, they score 41/100 and 33/100 respectively. Therefore, capabilities of LLMs for encoding real-world observational distributions seem limited, and using such knowledge for downstream tasks should be done with care.
Figure 4: Model leaderboard.
Furthermore, the analysis in Appendix LABEL:appendix:fine-tune shows that fine-tuning on samples from the ground truth dataset offers limited added benefit, raising interesting research questions on how to design approaches for improving the models’ L1 knowledge. In Appendix LABEL:appendix:fairness-impact we investigate the possible impact of debiasing techniques (tuning models for fairness and alignment), to understand if such methods may skew the results of our benchmark.
@@ 3 Results
Another interesting aspect is the performance of models across different datasets (low-dimensional setting), visualized in Fig. 5. The best performance is observed over the FBI Arrest Statistics (sex/race distributions over crime types), ACS Census data (questions on employment, education, and income across demographics), and Departments of Education (sex by graduation degree) and Labor (sex/race by occupation). For three of these four datasets (FBI, BLS, IPEDS) the detailed statistics we queried are available on their websites, meaning that models could have had access to the exact probability tables used for constructing our questions. Appendix LABEL:appendix:data-contamination investigates whether improved performance may correspond to the actual data being available in the model’s training set, but we find no relation between question entropy (a proxy for whether a question was seen during pre-training (hendrycks2020measuring)) and model performance, implying that question memorization may not play an important role in context of our benchmark.
@@
We next inspect model performance with varying dimension in the high-dimensional setting (Fig. 6(a)). The pattern of the performances does not clearly indicate the curse of dimensionality, possibly due to the fact that each dimension $d$ has a relatively small number of tasks, and additional tasks may need to be added to investigate this effect in more detail. However, overall, models perform better in the low-dimensional setting than in the high-dimensional setting (Fig. 4).
Figure 5: Performance by dataset (low-dimensional).
Closed and Reasoning Models.
We next compare the performance of closed models vs. open-weight models. In particular, we evaluate two closed models from OpenAI – GPT-4.1 (achiam2023gpt) and o4-mini (openai2025o4mini), where the latter model is a reasoning language model (RLM) (besta2025reasoninglanguagemodelsblueprint). Additionally, we also include a version of DeepSeek R1 (guo2025deepseek) distilled using Qwen2.5 32B (hui2024qwen2). As the OpenAI models require access to an API, we evaluate them on a subset of the tasks in the high-dimensional setting (consisting of tasks with a univariate binary $V_{Y}$, and $|V_{X}|\in\{2,3,4,5\}$). In particular, we select 63 out of 94 tasks (the selected tasks are indicated in the Closed Eval column of Tab. LABEL:tab:high-dim), focusing on tasks for which at most 250 queries need to be evaluated on the model. To elicit the model’s distribution $\tilde{P}(V_{Y}\mid V_{X})$, we use likelihood-prompting, since closed models do not provide access to next-token prediction probabilities, making QA prompting possible only via Monte Carlo sampling, which is prohibitively expensive.  [^1] Monte Carlo sampling would increase the number of queries sent to the models and the associated cost by at least two orders of magnitude, due to the fact that even with $n_{mc}=100$ Monte Carlo samples, the 95% confidence interval for a probability $\tilde{P}(V_{Y}=1\mid V_{X}=v_{X})\in[0,1]$ is only guaranteed to be reduced to a width of approximately 0.1, still reflecting a high level of uncertainty.
(a) Performance by dimension.
(b) Performance of GPT-4.1 and o4-mini.
Figure 6: Model performance for (a) dimension; (b) closed models.
The performance of GPT-4.1, o4-mini, and DeepSeek R1 (Qwen2.5 32B distill) models on the 63 selected tasks is shown in Fig. 6(b). The figure shows that GPT-4.1 and o4-mini outperform the evaluated open-weights models. This may be explained by the fact that GPT-4.1 is the largest model we evaluated, in terms of the parameter count (even though the number of parameters is not officially published). Even though OpenAI’s models outperform the evaluated open-weights models, they still do not perform well, which we further contextualize in the discussion of baselines below.
Retrieval-Augmented Generation (RAG) (lewis2020retrieval).
Retrieval-augmented generation (RAG) is a technique in which an LLM is combined with an external retriever (lewis2020retrieval). In this case, the model does not have to rely only on its internal knowledge, but relevant documents can be fetched from a collection of texts. To investigate the impact of RAG on our benchmark, we used the GPT 4.1 model with web access (the model can search the web for each query). To test how RAG affects performance, we focused on 11 tasks on which the GPT 4.1 model without web-based RAG had a score of 0 (corresponding to task indices 40, 43, 57, 62, 63, 64, 65, 66, 67, 70, and 71 in Tab. LABEL:tab:high-dim). Here, the expectation is that RAG would help the model performance, and would result in a higher score. However, the results for the GPT 4.1 model with web-based RAG still showed a zero score for each task. This finding warrants further investigation of how RAG may improve the performance of models on our benchmark.
Baselines – how good are the models?
In the evaluation of our benchmark, two baselines are used, corresponding to the notion of no probabilistic knowledge. The first baseline was the uniform baseline $P^{\text{unif}}$, which is in the binary case given by:
$P^{\text{unif}}(V_{Y}=1\mid V_{X}=v_{X})=\frac{1}{2}\;\forall v_{X}.$ (7)
This baseline corresponds to random guessing. For high-dimensional tasks with binary outcomes, we also used the 0/1 baseline, which outputs either a probability 0 for events with a marginal probability $\leq 0.5$, or a probability $1$ for events with a marginal probability >0.5>0.5:
where $\mathbbm{E}_{P}[V_{Y}]$ is the true mean of $V_{Y}$ (in the ground truth distribution $P$). In our scoring scheme, these baselines were used to determine the score $S=0$ for a given task (see Sec. 2.2.1). Clearly, the choice of baselines for determining the score $S=0$ has a strong effect on the evaluation. Here, another interesting baseline to consider is the mean-baseline, defined by:
This baseline outputs a fixed value (the marginal mean of the outcome $V_{Y}$) for each $v_{X}\in\mathrm{dom}(V_{X})$. In words, this baseline is aware of the marginal probability of the event $V_{Y}$, but has no knowledge about the variation of $\mathbbm{E}[V_{Y}\mid V_{X}=v_{X}]$ according to $v_{X}$ around the marginal mean $\mathbbm{E}_{P}[V_{Y}]$. This baseline corresponds to some probabilistic knowledge (marginal mean), but without any knowledge over how the probability varies across population subsets. In Fig. 6(b), the performance of the mean-baseline with respect to the current evaluation is shown (orange horizontal line), indicating an average score of 46/100. This means that the inclusion of this baseline would effectively render the scores of all models investigated in this manuscript close to 0. This observation once again confirms that the current generation of LLMs do not possess robust knowledge on real-world observational distributions, regardless of the strategy (QA or likelihood prompting) for eliciting the model distributions $\tilde{P}(V_{Y}\mid V_{X})$.
Summary of Supplementary Analyses.
In addition to the main results reported above, several supplementary analyses (Appendices LABEL:appendix:base-vs-it-LABEL:appendix:data-contamination) further contextualize the findings. Appendix LABEL:appendix:base-vs-it compares base and instruction-tuned variants of the same models, showing that instruction tuning does not substantially change models’ performance, in some cases leading to worsening and in some to improvement. This suggests that instruction-tuning is not a key aspect when analyzing L1 knowledge. Appendix LABEL:appendix:fine-tune examines the effect of fine-tuning on samples from the ground-truth dataset and shows that such fine-tuning offers only limited benefit for improving models’ L1 knowledge. Appendix LABEL:appendix:fairness-impact evaluates the impact of debiasing and alignment techniques on benchmark performance, assessing whether such interventions may systematically skew the elicited distributions. Appendix LABEL:appendix:data-contamination investigates potential data contamination by relating question entropy to model performance on different tasks, and finds no evidence that memorization meaningfully contributes to performance. Together, these analyses support the overall conclusion that current LLMs possess limited knowledge of real-world observational distributions.
4 Conclusion
In this work, we introduced the first benchmark designed to systematically evaluate whether LLMs can provide knowledge on observational distributions across diverse domains; or, in other words, whether their statistics are aligned with the world they are intended to model. Our benchmark assesses language models on a range of associational quantities for which the ground truth can be relatively well established using large-scale population surveys. The benchmark tests the hypothesis that LLMs can serve as a proxy for real-world correlations. However, results across a suite of models suggest that models’ statistics are consistently and substantially misaligned with the ground truth. We hope that these findings motivate research developments to improve the statistical fidelity of current language models.
@@ 4 Conclusion
As mentioned in the introduction, the recovery of observational distributions investigated in this work may be seen as related to questions of causal inference. Commonly, to adjust for confounding by a set of covariates $V_{Z}$, data analysts compute the propensity of a treatment $V_{X}$, that is, reweigh the samples with $P(V_{X}=v_{X}\mid V_{Z}=v_{Z})^{-1}$ to obtain estimates of a desired potential outcome (rubin1977assignment; rosenbaum1983central). Other strategies use a predictive model for the outcome for the outcome distribution $P(V_{Y}\mid V_{X},V_{Z})$ in order to allow for or improve causal inference (bang2005doubly). However, if such conditional distributions are not available to a model, eliciting them may be a futile exercise (see Cor. 2). Therefore, our work brings into question some recent approaches claiming to improve causal inferences using foundation models (de2025efficient). In our view, using AI models for causal inference may require a more sophisticated approach.
@@
We acknowledge some limitations of our work. Currently, a limited number of models are evaluated on our benchmark. However, by choosing a representative sample of models, we cover a range of models developed by different organizations. Furthermore, our emphasis on easy access and evaluation of models will facilitate the benchmarking of other models in future work. Another limitation is that high-dimensional inference is investigated across 4 datasets, and the number of contexts for evaluation should grow over time. Additionally, we mention that all datasets considered describe US populations, and future work should aim for wider geographic coverage. We further envision that our work will set the basis for evaluating capabilities of AI models for inference in Layers 2 & 3 of the PCH – to systematically test how observational distributions available to models affect downstream tasks such as causal or counterfactual inferences.
@@
The source code for reproducing the benchmark can be found in our code repository https://github.com/dplecko/llm-epidemia. The repository includes a README file explaining all the steps for setting up the benchmark. Our Hugging Face repository https://huggingface.co/spaces/llm-observatory/llm-observatory-eval includes code for streamlined evaluations of models on the benchmark. Benchmark results can also be inspected on the project webpage https://llm-observatory.org/. For the Llama3 8B, Mistral 7B, and Phi4 models the experiments were run on a single NVIDIA H100 GPU. Evaluating both the low- and high-dimensional settings required <1 hour per model. Evaluating DeepSeek 7B (Chat version), Gemma3 27B, and LLama3 70B models was done on a node with four NVIDIA GH200 Grace Hopper chips, and running both settings required <1 hour of compute per model.
@@ 74 SCF: Debt by Education
Table 2: High-dimensional tasks. The third column indicates the task dimension $d=|V_{X}|$, and the fourth column indicates whether the task is used for the evaluation of closed models in Fig. 6(b).
@@ Appendix B Comparison of Base & Instruction-Tuned Models
Appendix C Does Fine-Tuning Improve Performance?